Blossom Company management is considering a project that will require an initial investment of $45,000 and will last for 10 years. No other capital expenditures or increases in working capital are anticipated during the life of the project. What is the annual EBIT that will make the project economically viable if the cost of capital for the project is 8 percent and the firm will depreciate the investment using straight-line depreciation and a salvage value of $0? Assume that the marginal tax rate is 30 percent. (Do not round intermediate calculations. Round final answer to 2 decimal places, e.g. 52.75.)


Please show how to use the 8% in sum of factors calc
Answer #1

NPV = cash outflow + net operating cash flow*PVAF at 8% for 10 years

0 =-45000+net operating cash flow*PVAF

0=-45000+operating cash flow*6.7100

0= -45000+6.7100 operating cash flow

Operating cash flow = 45000/6.7100 = 6706.327

PVAF at 8% for 10 Years = 1-(1+r)^-n / r = 1-(1.08)^-10 / .08 = 6.7100

EBIT = 2206.4/70%

3152

less tax - 30% of EBIT

3152*30%

945.6

PAT-70% 0f EBIT

6706.4-4500

2206.4

depreciation

4500

operating cash flow

6706.4

EBIT at which proposal is beneficial at 8%

3152